Rod is rotated with constant angular velocity omega about the vertical axis. Find tension in the rod at a distance x from the axis of rotation.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
AB is a rod of length l ,mass M & this rod is rotating about end A....
now , mark a point c anywhere on the rod , distance bw end A & C is x ...
now the rod is divided into two parts , AC & CB .......
tension of rod at point C is (T)= dmw2r                   (dm is mass of part BC = m(l-x)/l )
                                          =M(l-x)w2r/l
this r is the distance of center of mass of BC from end A so r = (x+l)/2
so
           T = M(l-x)(l+x))w2 /2l    
          T = M(l2-x2)w2 /2l